how do I go about solving linear equations: -x-47=4 and x+4y=4. I need to state no solution, unique or infinity? I was trying to use calculator but it wouldn't let me enter a -x value.
Answer by stanbon(75887)

how do I go about solving linear equations: -x-4y=4 and x+4y=4. I need to state no solution, unique or infinity?
------
Add the two equations to get:
0+0 = 8
-----
That is a contradiction.
So the system has no solution.
----
Cheers,
Stan H.
---------